OpenStudy (anonymous):

Is there a rule to how long you should wait until reposting something? I would really like to get this question out of the way ^^;

OpenStudy (amistre64):

one rule says 3 days :)

OpenStudy (anonymous):

If you can't see it in a full scroll, you can go ahead and re-post Particularly if you are going to be around to work with a tutor immediately.

OpenStudy (anonymous):

Alright, thanks.

OpenStudy (anonymous):

Is this the question? Pistachios cost $6.50/pound and Almonds cost $8.00/pound. How much of each to make a 50lb mixture that costs $7.50/pound? Any help in the right direction would be great!

OpenStudy (anonymous):

Yes, it is.

OpenStudy (anonymous):

Ok well Lets use P for the amount of pistachios and A for the amount of almonds.

OpenStudy (anonymous):

So the resulting 50 pounds relates to P and A how?

OpenStudy (amistre64):

ill abstract it just so as not to stir the kettle to bad lol a+b = t; a = t-b ax + by = tz (t-b)x + by = tz tx -bx + by = tz -bx + by = tz - tx b(y-x) = t(z-x) b@y = t(z-x)/(y-x) ; :)

OpenStudy (anonymous):

lol, I know how to do this and even I can't make sense of that amistre ;p

OpenStudy (anonymous):

I believe something like: 6.50p + 8.00a = 50

OpenStudy (anonymous):

ooh, good guess

OpenStudy (anonymous):

But look at the unit analysis.

OpenStudy (anonymous):

Not sure what unit analysis is. ^^;

OpenStudy (anonymous):

6.5(dollars/pound) * p (pounds) + 8(dollars/pound) * a (pounds) = dollars But you equated them to 50 pounds

OpenStudy (amistre64):

another method I got brewing in me head is: a+ b = t ; multiply by say, -x xa + yb = tz -xa - xb = -tx xa + yb = tz -------------- b(y-x) = t(z-x) ... lol, it comes out the same

OpenStudy (anonymous):

Oh, I see. So we need: 6.50p + 8.00a = 7.40?

OpenStudy (anonymous):

but 7.5 is in dollars/pound not just dollars

OpenStudy (anonymous):

so how many dollars will 50 pounds of the mixture be if it's 7.5 dollars per pound?

OpenStudy (anonymous):

370

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

So now we can solve for the amount of one in terms of the other.

OpenStudy (anonymous):

But we still need to know something more.

OpenStudy (anonymous):

How so? It's 7.40 per pound, not 7.50

OpenStudy (anonymous):

"How much of each to make a 50lb mixture that costs $7.50/pound?"

OpenStudy (anonymous):

Oh! I'm sorry, I typo'd the first post. It's 7.40 per not 7.50!

OpenStudy (anonymous):

oh, well then make the appropriate adjustment.

OpenStudy (anonymous):

So we have now: 6.50p + 8.00a = 370?

OpenStudy (anonymous):

So would I use the subsitution method to solve for one of the variables?

OpenStudy (anonymous):

Well yes, but we need one more equation to be able to substitute.

OpenStudy (anonymous):

That's true, lol.

OpenStudy (anonymous):

We have 1 equation and 2 unknowns. But let me ask you this. That 50 pounds we were talking about earlier. How does it relate to P and A?

OpenStudy (anonymous):

Because the 50 pounds is a mix of P (Pistachios) and A (Almonds) to equal 50lbs?

OpenStudy (anonymous):

Right. so can we use that to write another equation?

OpenStudy (anonymous):

P is in units of?

OpenStudy (anonymous):

We said that P was the 'amount' of pistachos, but really what do we mean by 'amount'?

OpenStudy (anonymous):

By amount we mean pounds I believe. So we have x pounds of P.

OpenStudy (anonymous):

No, we have P pounds of pistachios

OpenStudy (anonymous):

Ah, hmm, lets see.

OpenStudy (anonymous):

But yes, it's in pounds. So if we take the pounds of pistachios and the pounds of almonds what is the relation to the final weight of the mixture?

OpenStudy (anonymous):

AllI can think of is that they have the relation that together they = 50lb but we already knew that.

OpenStudy (anonymous):

Quick question: Would we want to take the decimals out of this equation?

OpenStudy (anonymous):

I apologize for the abrupt leave but I must go. I hope we can pick this up at a later time. Thank you for your help so far. Glad I found this place!

OpenStudy (anonymous):

AllI can think of is that they have the relation that together they = 50lb but we already knew that. We do know this, but we haven't used it yet! This is the second equation we need: P + A = 50

OpenStudy (anonymous):

As for the decimals, you can keep them or just multiply the whole equation by 10 it amounts to the same thing.

OpenStudy (anonymous):

Sorry for not getting back to you sooner, I had a phone call. Hopefully you'll see this later.